If you had excess chlorine, how many moles of of aluminum chloride could be produced from 23.0 g

g of aluminum?


1
Expert's answer
2021-11-22T00:18:02-0500

From balanced chemical equation we see that 2 mol Al forms 2 mol AlCl3


0.667 mol Al will forms = 0.667 mol AlCl3 Ans.


moles of Cl2 = 23.0g /70.906 g/mol = 0.324 mol


3 mol Cl2 gives 2 mol AlCl3


0.324 mol will give = 2/3 x 0.324 = 0.216 mol AlCl3